Use the properties of summation and the Summation Formulas Theorem to evaluate the sum. Use the summation capabilities of a graphing utility to verify your result.

18 (i − 1)2
i = 1
Answer #1


is g. for 5 [ (ie 5 ( 12+1-21) = i=1 +51-22 (1) = 1 4-1 uring formulas 18(18+)(241871) + 18 - 21584E+) (1785 |